A linear regression model predicting the linear effects of the amount of water used in growing almonds is developed below:

y = 56+10x

The water data has a range of data from 5 to 90 and is in units of 100 gallons, and the almond data is in pounds.

According to the model, how many acres will grow if 1000 gallons is used
Answer #1

1000 gallons means 10 ( 100 gallons )

y = 56 + 10x

y = 56 + 10*10

y = 56 + 100

y = 156

almonds = 156 pounds